# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v11

@@ 8.7 Using Library Sine for Sin/Cos Signals
Figure 37: FFT spectrum of $\sin(2\pi 3j/2^{6})$ using Library sine function after the forward transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 38: FFT waveform of $\sin(2\pi 3j/2^{6})$ using Library sine function after the reverse transformation computed by variance arithmetic. The legend distinguishes between the uncertainty and the value errors. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 39: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin signal at frequency 33 as a function of FFT order (x axis) and transformations (legend) using Library sine function.
Figure 40: Histograms of normalized errors of Sin signal at frequency 33 for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
Using library sine function, for a sine wave with $3$ as frequency, Figure 37 displays the spectrum for the forward transformation, while Figure 38 presents the waveform for the reverse transformation:
  • The uncertainties are identical to the corresponding values using Quart sine function, because the input uncertainties are identical for these two sine functions. This identical relationship presents in all FFT orders when comparing the corresponding uncertainty means in Figure 35 and 39.
  • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in Library sine function are more than those using Quart sine function in Figure 31.
When compared with Figure 36, Figure 40 also confirms that at FFT order $18$, normalized errors have much wider distributions using Library sine function than those using Quart sine function.
Due to numerical errors in Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 39.
@@ 8.8 Numerical Error Resonance Using Library Sine
Figure 41: Error deviation (z-axis) of FFT forward transformation of $sin(2\pi fj/2^{L})$ as a function of frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
Figure 42: Error deviation (z-axis) of FFT reverse transformation of $sin(2\pi fj/2^{L})$ as a function of frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
Value errors in the reverse transformation show a clear trend of increasing with time index, as in Figure 41. These large value errors looks systematic and do not appear as typical noise visually; or rather, they look more like a resonant pattern. Such increases present in other frequencies and FFT order, as well as in computational result using mathematical libraries such as SciPy. A FFT spectrum of the forward transformation shows a sudden increase of intensity at frequency index which is also consistent with the result of SciPy, as in Figure 37. Figure 41 and 42 show that error deviations of such value errors increase with sine or cosine frequencies, the resonant beats between the signal and the numerical library errors become stronger. In contrast, such increase does not appear at all when using Quart sine function, such as in Figure 33 and 34. Figure 31 shows that the numerical errors using Library sine increase with a periodicity of $\pi$, which may resonate with a signal which has a periodicity of an integer fold of $\pi$, to produce the resonant pattern in Figure 37 and 38. To avoid the numerical error resonance in Figure 41 and 42, $10^{-14}$ input noise is required to add to the input sine or cosine signals.
@@ 8.9 Using Quart Sine for Linear Signal
Figure 43: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal as a function of FFT order (x axis) and transformations (legend) using Quart sine function.
Figure 44: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) using Quart sine function. The FFT order is $18$.
Figure 43 shows that using Quart sine function can track the result uncertainties of Linear signals with proper coverage. Because the input to reverse transform no longer contains mostly precise zeros, the output uncertainty increases far more rapidly with FFT order than its counterpart in Figure 35.
Figure 44 shows that for each transformation, the normalized error distribution in reverse transformation is similar to its counterpart of the Sin/Cos signals using Quart sine function in Figure 36, while the distribution in the forward transformation is similar to the Linear signal using Library sine function in Figure 46. The reason for this disparity in similarity of the normalized error distributions is not clear.
@@ 8.10 Using Library Sine for Linear Signal
Figure 45: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal as a function of FFT order (x axis) and transformations (legend) using Library sine function.
Figure 46: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
As shown as the difference between Figure 31 and 32, linear signals using Library sine function may introduce $10^{3}$ times larger numerical errors into the results through Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these extra errors or not.
Figure 45 shows that proper coverage can no longer be achieved because value error outpaces uncertainty as FFT order increases. Figure 46 shows that the normalized error distribution for the reverse transformation is no longer bound effectively within three deviations. Although the normalized error distribution for the forward transformation using Library sine function looks similar to its counterpart of using Quart sine function in Figure 46 looks similar to that using Quart sine function in Figure 44, detailed analysis shows that the normalized error distribution using Library sine function contains extreme values at its two long tails, such as $2.8\;10^{3}$ and $3.9\;10^{3}$ as the minimal and maximal, respectively. The presence of these extreme values consistent with those delta-like numerical errors from Figure 32.
The observed difficulty of variance arithmetic in tracking Linear signals using Library sine function suggests that variance arithmetic may fail when the input contains excessive unspecified errors. A potential solution is to develop a new sin library implemented within variance arithmetic framework, ensuring that all numerical calculation errors are explicitly accounted for, as what Quart sine function has achieved in FFT transformations.
@@ 8.11 Ideal Coverage
Figure 47: Histograms of normalized errors of Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) using Library sine function. The FFT order is $18$.
Figure 48: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals as a function of FFT order (x axis) and transformations (legend) using Library sine function.
Figure 49: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for forward transformations of Linear signals using Library sine function.
Figure 50: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Library sine function.
Figure 51: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for forward transformations of Linear signals using Quart sine function.
Figure 52: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Quart sine function.
Adding sufficient noise to the input can overpower unspecified input errors, thereby achieving ideal coverage; for example, applying an input noise of $10^{-3}$ to a Linear signal when using library sine function.
  • Figure 47 presents the corresponding histogram. As expected, the normalized errors for the forward and reverse transformations follow Normal distributed, while those for the roundtrip transformations are Delta distributed at 0zero, indicating perfect recovery of the input uncertainties.
  • Figure 48 illustrates the corresponding error deviations and uncertainty means:
@@
    • As expected, the result error deviations for the forward and reverse transformations are constant at $1$, whereas those for the roundtrip transformation decay exponentially to 0 with increasing FFT order.
@@
Additionally, the result uncertainty means for both forward and reverse transformations are linearly proportional to the input uncertainties, which is expected because FFT is a linear algorithm [13].
@@
The range of ideal coverage depends on how accurately the input uncertainty reflects the actual input noise. For Linear signals using Library sine function, Figure 49 and 50 present the error deviation as a function of the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation is $1$. Outside this region, proper coverage cannot be achievable. Because uncertainties grow more slowly in reverse transformation than in forward transformation, the reverse transformations exhibit a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that yields ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually, beyond FFT order $25$ for the reverse transformation, ideal coverage may not be achievable. Although FFT is regarded as one of the most robust numerical algorithms very insensitive to input errors, it can still fail due to numerical errors in library sine function. Such deterioration of the calculation is not easily detectable using conventional floating-point arithmetic.
@@
In contrast, for Linear signals using Quart sine functions, as shown in Figure 51 and 52 for the forward and reverse transformations, respectively, the ideal coverage region is significantly larger, and proper coverage is also achieved in other regions. Forward transformations show a larger ideal coverage region than that of reverse transformations.
@@
As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders. The difference is that using Quart sine functions, within the proper coverage region, error deviations differ from $1$ only marginally.
8.12 Prec Sine Functions
@@
Figure 53: Value Errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) as a function of $x$ (x-axis) and $L$ using either Prec or Quart sine function (legend).
@@
Figure 54: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals as a function of FFT order (x axis) and transformations (legend) using Prec sine function.
@@
Figure 55: Histograms of normalized errorsfor forward, reverse and roundtrip transformations (legend) of Linear signals using Prec sine functions.
@@
Figure 56: Error deviations (z-axis) as a function of input uncertainties (x-axis) and FFT orders (y-axis) for reverse transformations of Linear signals using Prec sine function.
@@
Quart sine function over-estimates the uncertainty, while Library sine function underestimates the uncertainty. Is it possible to have a better sine function? If Quart sine function is constructed using the 128-bit floating-point library of gcc, then converted to 64-bit values, using the difference between a 128-bit and the corresponding 64-bit value as the uncertainty for the 64-bit value, the resulting indexed sine function is named as the Prec sine function.
@@
Using $\sin(x)^{2}+\cos(x)^{2}-1$ test to compare Prec and Quart sine functions, Figure 53 illustrates that Prec sine function produces slightly more value errors, while having half uncertainties. Figure 54 and 55 both show that Prec sine function underestimates uncertainties slight, with stable error deviation around $1.8$ for both forward and reverse transformations without added noise. Compared with 52, Figure 56 illustrates that the ideal coverage using Prec sine function is identical to using Quart sine function. The observation that 128-bit sine function does not improve over 64-bit sine function in any noticeable way challenges the legitimacy of using the difference as the uncertainty, the correctness of the 128-bit gcc numerical library, or the algorithm in converting a 128-bit floating-point value to 64-bit.
@@
Both Quart and Prec sine function are not ideal. An ideal sine function should have error deviations closer to $1$ without added noise.
@@ 8.13 Summary
The library sine functions implemented with conventional floating-point arithmetic have been shown to contain numerical errors equivalent to $10^{-3}$ of input precision in worst cases for FFT transformations. These library $\sin(x)$ errors increase periodically with $x$, which may resonate with periodic input signals. The computational result using floating-point arithmetic is very sensitive to input data, amount of calculation, and library function implementation. This means that a small-scale test cannot properly qualify the result of a large-scale calculation. The impact of numerical errors within mathematical libraries has not received sufficient attention. The numerical errors using floating-point arithmetic are very difficult to detect and quantify.
Variance arithmetic offers a robust alternative, as its computed values closely match those from conventional floating-point arithmetic while its associated uncertainties trace all input errors. Moreover, its resulting error deviations enable objective classification of calculation quality as ideal, proper, or suspicious.

### v12

@@ 9.7 Using Library Sine for Sin/Cos Signals
Figure 40: FFT spectrum of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the forward transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the frequency index, while the y-axis represents both the uncertainty and the value error.
Figure 41: FFT waveform of $\sin(3\frac{2\pi}{2^{6}}j)$ computed using either the Library sine function or SciPy after the reverse transformation. The legend distinguishes between the uncertainty and the value error. The x-axis represents the time index, while the y-axis represents both the uncertainty and the value error.
Figure 42: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Sin/Cos signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
Figure 43: Histograms of normalized errors of Sin/Cos signal for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
Using the Library sine function, for a sine wave with a frequency of $3$, Figure 40 displays the spectrum for the forward transformation, while Figure 41 presents the waveform for the reverse transformation:
  • The uncertainties are identical to the corresponding values obtained using the Quart sine function, since the input uncertainties are the same for both sine functions. This identical relationship persists across all FFT orders when comparing the corresponding uncertainty means in Figure 38 and 42.
  • The error deviations for the forward and reverse transformations are $3.7$ and $3.0$ respectively, confirming that the numerical errors in the Library sine function are greater than those using the Quart sine function, as illustrated in Figure 34.
  • When compared with Figure 39, Figure 43 also confirms that at FFT order $18$, the normalized errors distribution is significantly broader when using the Quart sine function.
  • Due to numerical errors in the Library sine function, error deviations reach their stable values much quicker with FFT order in Figure 42 that that in Figure 38.
Using the Library sine function, variance arithmetic also achieves proper coverage for sine and cosine signals.
@@ 9.8 Numerical Error Resonance Using Library Sine
Figure 44: Error deviation (z-axis) of FFT forward transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
Figure 45: Error deviation (z-axis) of FFT reverse transformation of $sin(f\frac{2\pi}{2^{L}}j)$ versus frequency $f$ (x-axis) and FFT Order $L$ (y-axis).
In the reverse transformation, value errors exhibit a clear trend of increasing with the time index, as shown in Figure 44. These large value errors appear systematic rather than random and visually resemble a resonant pattern. Similar increases are observed at in other frequencies and FFT orders, as well as in computational results obtained using mathematical libraries such as SciPy. In the forward transformation, although there is no visible resonant pattern in FFT spectrum, the intensity of value errors increase with FFT order. Figure 44 and 45 demonstrate that the error deviations increase with sine or cosine frequency. In contrast, this increase is completely absent when using the Quart sine function, as seen in Figure 36 and 37. Figure 34 indicates that the numerical errors using the Library sine function increase with a periodicity of $\pi$, which may resonate with a signal whose periodicity of an integer multiply of $\pi$, producing the resonant pattern in Figure 40 and 41. At higher frequency, the resonant beats between the signal and the numerical errors in the Library sine function become stronger. To suppress this numerical error resonances, an input noise level of approximately $10^{-14}$ input noise must be added to the sine or cosine signals.
@@ 9.9 Using Quart Sine for Linear Signal
Figure 46: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Quart sine function.
Figure 47: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Quart sine function. The FFT order is $18$.
Figure 46 shows that using the Quart sine function, the result uncertainties can track the value errors of Linear signals with proper coverage. Because the input to the reverse transformation no longer consists predominantly of precise zeros, the output uncertainty increases much more rapidly with FFT order than its counterpart in Figure 38. This increase enables the proper coverage in Figure 46, resulting in error deviations similar to those Figure 38. The normalized error histogram of the reverse transformation in Figure 49 is narrower than that of Figure 43, confirming Figure 46. Using the Quart sine function, variance arithmetic can provide proper coverage for Sin signals, Cos signals, and Linear signal.
@@ 9.10 Using Library Sine for Linear Signal
Figure 48: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signal versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
Figure 49: Histograms of normalized errors of Linear signals for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
As shown by the difference between Figure 34 and 35, Linear signals computed using the Library sine function can introduce unspecified numerical errors up to $10^{3}$ times larger into the results through the Library $\cos(x)/sin(x)$. The key question is whether variance arithmetic can effectively track these additional errors.
Figure 48 shows that proper coverage can no longer be achieved, as the value error increases faster than the uncertainty with rising FFT order. Figure 49 demonstrates that the normalized error distribution for the reverse transformation is no longer effectively bounded within three standard deviations. Although the normalized error distribution for the forward transformation using the Library sine function appears visually similar to that obtained using the Quart sine function in Figure 49, a detailed analysis reveals that the normalized error distribution contains extreme values at its two long tails, such as $2.8\;10^{3}$ and $3.9\;10^{3}$ as the minimal and maximal of the normalized errors, respectively. The presence of these extreme values is consistent with those delta-like numerical errors observed in Figure 35.
The observed difficulty of variance arithmetic in tracking Linear signals computed using the Library sine function suggests that variance arithmetic may fail when the input contains excessive unspecified numerical errors. A potential solution is to develop a new sine function library implemented directly within the variance arithmetic framework, ensuring that all numerical calculations are performed with explicit uncertainty tracking.
@@ 9.11 Ideal Coverage
Figure 50: Result error deviation (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT order (x-axis) and transformation types (legend) computed using the Library sine function.
Figure 51: Histograms of normalized errors for Linear signals with $10^{-3}$ input noise for forward, reverse and roundtrip transformations (legend) computed using the Library sine function. The FFT order is $18$.
Figure 52: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Library sine function.
Figure 53: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Library sine function.
Figure 54: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the forward transformations of Linear signals computed using the Quart sine function.
Figure 55: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for the reverse transformations of Linear signals computed using the Quart sine function.
Adding sufficient noise to the input can suppress unspecified input errors and thereby achieve ideal coverage. For example, applying an input noise of $10^{-3}$ to a Linear signal when using the Library sine function result in ideal coverage.
Figure 50 illustrates the error deviations and uncertainty means:
@@
  • As expected, the result error deviations for the forward and reverse transformations remain constant at $1$, whereas those for the roundtrip transformation decay exponentially toward 0 as the FFT order increases.
Figure 51 presents the corresponding histogram. As expected, the normalized errors for the forward and reverse transformations follow Normal distributions, while those for the roundtrip transformation are Delta distributed around zero, indicating perfect recovery of the input uncertainties. According to Figure 51, adding either Gaussian or white noise to input results in the same Gaussian distribution of the normalized errors.
@@
Additionally, the result uncertainty means for both forward and reverse transformations are linearly proportional to the input uncertainties, as expected since FFT is a linear algorithm [13].
@@
The range of ideal coverage depends on how accurately the specified input uncertainty represents the actual input noise. For Linear signals computed using the Library sine function, Figure 52 and 53 present the error deviation versus the added noise and FFT order for the forward and reverse transformations, respectively. Ideal coverage corresponds to the region where the error deviation equals $1$. Outside this region, proper coverage cannot be achieved. Because uncertainties grow more slowly in the reverse transformation than in the forward transformation, the reverse transformation exhibits a smaller ideal coverage region. Furthermore, as numerical errors increase with computational load, the range of input noise that produces ideal coverage decreases with increasing FFT order. At sufficiently high FFT orders, visually beyond FFT order $25$ for the reverse transformation, ideal coverage may no longer be achievable. Although FFT is widely regarded as one of the most robust numerical algorithms, and generally insensitive to input errors, it can still fail due to numerical errors in the Library sine function. Such deterioration in calculation accuracy is not easily detectable when using conventional floating-point arithmetic.
@@
In contrast, for Linear signals computed using the Quart sine functions, Figure 54 and 55 show the results for the forward and reverse transformations, respectively. The ideal coverage region is significantly larger, and proper coverage is also achieved in adjacent regions. The forward transformations exhibits a broader ideal coverage region than the reverse transformations.
@@
As a comparison, because Sin/Cos signals have fewer numerical calculation errors, using either Quart or the Library sine functions, the ideal coverage region is achieved once the added noise is large enough to cover the effect of rounding errors, and this condition is almost independent of FFT orders. The key difference is that when using the Quart sine functions, the error deviations within the proper coverage region differ from $1$ only marginally.
9.12 Prec Sine Function
@@
Figure 56: Value errors of $\sin(x)^{2}+\cos(x)^{2}-1,x=\pi j/2^{L}$ (y-axis) versus $x$ (x-axis) and $L$ computed using either the Prec or the Quart sine function (legend).
@@
Figure 57: Error deviations (left y-axis) and uncertainty mean (right y-axis) of Linear signals versus FFT orders (x-axis) and transformation types (legend) computed using the the Prec sine function.
@@
Figure 58: Histograms of normalized errorsfor forward, reverse and roundtrip transformations (legend) of Linear signals computed using the the Prec sine function.
@@
Figure 59: Error deviation (z-axis) versus input uncertainty (x-axis) and FFT order (y-axis) for reverse transformations of Linear signals computed using the the Prec sine function.
Signal vs Sine Transformation Quart Prec Library
Sin/Cos Forward $10^{-15}$ $10^{-15}$ $10^{-14}$
Sin/Cos Reverse $10^{-12}$ $10^{-12}$ $10^{-11}$
Linear Forward $10^{-9}$ $10^{-9}$ $10^{-6}$
Linear Reverse $10^{-7}$ $10^{-7}$ $10^{-3}$
Table 2: The minimal required noise to achieve ideal coverage for FFT transformations at FFT order $18$ for different sine functions.
@@
The Quart sine function tends to overestimate the uncertainty, while the Library sine function underestimates it. This raises the question of whether a sine function with a more accurate uncertainty can be developed. If the Quart sine function is constructed using the 128-bit floating-point library of gcc and then converted to 64-bit values, the difference between each 128-bit value and its corresponding 64-bit value can be used as the uncertainty for the 64-bit representation. The resulting indexed sine function is referred to as the Prec sine function.
@@
Using $\sin(x)^{2}+\cos(x)^{2}-1$ test to compare the Prec and the Quart sine functions, Figure 56 shows that the the Prec sine function produces slightly larger value errors, while having half uncertainties. Figure 57 and 58 demonstrate that the the Prec sine function slightly underestimates uncertainties, maintaining a stable error deviation around $1.8$ for both forward and reverse transformations without added noise. Compared with Figure 55, Figure 59 shows that the ideal coverage achieved using the the Prec sine function is identical to that obtained with the Quart sine function. The observation that the 128-bit sine function provides no noticeable improvement over the 64-bit sine function raises concerns about the statistical legitimacy of directly using the difference as the uncertainty, the accuracy of the gcc 128-bit numerical library, or the algorithm used to convert a 128-bit floating-point values to 64-bit.
@@
Table 2 compares using different sine functions for Sin/Cos or Linear signals. Both Quart and the the Prec sine function are not ideal. An ideal sine function should have error deviations closer to $1$ without added noise.
@@ 9.13 Summary
The Library sine functions implemented using conventional floating-point arithmetic have been shown to contain numerical errors equivalent up to $10^{-3}$ of the input precision in the worst cases of FFT transformations. These $\sin(x)$ numerical errors increase periodically with $x$, potentially resonating with periodic input signals. Computational result obtained with floating-point arithmetic are highly sensitive to the input data, the scale of computation, and the specific implementation of the mathematical library. This implies that a small-scale test cannot reliably qualify or predict the behavior of large-scale calculations. The impact of numerical errors within mathematical libraries has not received sufficient attention. Moreover, such errors in floating-point arithmetic are inherently difficult to detect and quantify.
Variance arithmetic offers a robust alternative, as its computed values closely align with those obtained from conventional floating-point arithmetic while its associated uncertainties systematically trace all input errors. Furthermore, its resulting error deviations provide an objective basis for classifying of calculation quality as ideal, proper, or suspicious.